OpenStudy (anonymous):

Solve the quadratic equation by isolating the perfect square. 2(x+4)^2-2=34 I got x=-1 but my teacher marked that wrong?

OpenStudy (anonymous):

did you try expanding it and factoring it?

OpenStudy (anonymous):

I added 2 to both side. 2(x+4)^2 = 36 I then square rooted both sides. 2(x+4) = 6 I divided by two. x+4=3 I subtracted 4. x=-1

OpenStudy (anonymous):

you should divide both sides by two first before you get the square root if you get the square root of a number it should be positive or negative

OpenStudy (anonymous):

what happened to GEMDAS

OpenStudy (anonymous):

grouping first i thought

OpenStudy (anonymous):

just think it like this.. if you get the square root of both sides: